# Refined asymptotics for the Cauchy problem for the fast $p$-Laplace evolution equation

### v2

@@ 10 Comments, Extensions, and Open Problems
Lastly, let us comment on two very natural directions that may arise after the present work: the doubly nonlinear equation and anisotropic $p$-Laplace evolution equation. By the doubly nonlinear diffusion equation we mean $\partial_{t}u=\Delta_{p}(u^{m})$. The fast diffusion regime is when $p(m-1)<1$. It is known, at least in the corresponding good diffusion range, that (non-negative and integrable) solutions to the Cauchy problem behave for large times as the corresponding Barenblatt profiles, see for instance [1, 2, 3]. Of course, the very natural question is how much of what has been proven in this work also applies to doubly nonlinear case. We believe that the available regularity theory, see for instance [11, 12, 13, 67], allows to try to address questions (Q-1), (Q-2), and (Q-3).
@@ A.1 Lemmata
Let $21<p<$ 2 and $v:{\mathbb{R}^{N}}\rightarrow\left[0,\infty\right)$ be a measurable function. Suppose that there exists $D>0$ such that
@@
$\begin{split}0\leq V_{D_{2}}(y)-V_{D_{1}}(y)&=\int_{D_{1}}^{D_{2}}\frac{\partial}{\partial D}V_{D}(y)\,{\rm d}D=-\frac{p-1}{2-p}\int_{D_{1}}^{D_{2}}V_{D}^{\frac{1}{p-1}}(y)\,{\rm d}D\leq\frac{p-1}{2-p}|D_{2}-D_{1}|\,V_{D_{1}}^{\frac{1}{p-1}}(y)\,,\end{split}$
@@
Suppose $21<p<$ 2. Then there exist $c_{1},c_{2}>0$ such that for all $\xi,\eta\in{\mathbb{R}^{N}}$ such that $\xi\neq 0$ we have
@@ A.2 Parameters
$R_{T}(t)=\begin{cases}\{(T-t)_{+}/|\beta|\}^{\beta}\,,&1<p<p_{c}\,,\\ \exp\{\ell(T+t)\}\,,&p=p_{c}\,,\\ \{(T+t)/|\beta|\}^{\beta}\,,&p_{c}<p<2\,.\\ \end{cases}$

### v3

@@ 10 Comments, Extensions, and Open Problems
Lastly, let us comment on two very natural directions that may arise after the present work: the doubly nonlinear equation and anisotropic $p$-Laplace evolution equation. By the doubly nonlinear diffusion equation we mean $\partial_{t}u=\Delta_{p}(u^{m})$. The fast diffusion regime is when p(m1)<1p(m-1)<1. It is known, at least in the corresponding good diffusion range, that (non-negative and integrable) solutions to the Cauchy problem behave for large times as the corresponding Barenblatt profiles, see for instance [1, 2, 3]. Of course, the very natural question is how much of what has been proven in this work also applies to doubly nonlinear case. We believe that the available regularity theory, see for instance [11, 12, 13, 68], allows to try to address questions (Q-1), (Q-2), and (Q-3).
@@ Acknowledgements
The authors are deeply grateful to Professor Jean Dolbeault for bringing to their attention the reference [33], which had not been included in earlier versions of this manuscript.
@@ A.1 Lemmata
Let $1<p<2$ and $v:{\mathbb{R}^{N}}\rightarrow\left[0,\infty\right)$ be a measurable function. Suppose that there exists $D>0$ such that
@@
[…]
@@
Suppose $1<p<2$. Then there exist $c_{1},c_{2}>0$ such that for all $\xi,\eta\in{\mathbb{R}^{N}}$ such that $\xi\neq 0$ we have
@@ A.2 Parameters
[…]